OpenStudy (anonymous):

write ans in a positive exponent x^-5/x^-9

OpenStudy (anonymous):

switch them x^(-1) = 1/x

OpenStudy (anonymous):

this is a math law

OpenStudy (anonymous):

so 1/x^(-1) = x because 1/1/x = x

OpenStudy (anonymous):

now you should be able to solve this question unless you don't comprehend

OpenStudy (anonymous):

i know u have cant divide with a neg exponent but hows that ans make sense? dont get what u wrote sorry

OpenStudy (anonymous):

so basically ur saying its x^-5/1/x^9?

OpenStudy (anonymous):

No x^(-1) = 1/x so x^(-2) = 1/x^(2) That is a rule any negative exponent results in the number being transfered to the denominator. Knowing this rule we now know that x^-5 = 1/x^(5) x^-9 = 1/x^(9)

OpenStudy (anonymous):

ok... so far i understand

OpenStudy (anonymous):

1/x^(5)/1/x^(9) do you know how to divide fractions?

OpenStudy (anonymous):

well?

OpenStudy (anonymous):

sorry ya subtract the exponents

OpenStudy (anonymous):

noooo

OpenStudy (anonymous):

(1/2) / (1/3) = (1/2) * (3/1)

OpenStudy (anonymous):

ok sorry its been a long time since i did math sorry

OpenStudy (anonymous):

don't appologize you are still learning :)

OpenStudy (anonymous):

do you understand how to solve this question now or do you want me to show you the next step

OpenStudy (anonymous):

sort of need the help still... :-((

OpenStudy (anonymous):

ahhh i think i got it ..... i have to multiply 1/x^5 * x^9/1

OpenStudy (anonymous):

Whenever dividing two fractions (remember all numbers are fractions, whole numbers are just divide by 1, so 3 = (3/1) we just dont show it) to divide fractions you flip the lower number and multiply them so (x/2) / (x/3) = (1/2) * (3/x) = 1(3)/x(2) = 3/2x

OpenStudy (anonymous):

yup :) and the rules for multiply fractions is as follows (x/z)(y/a) = xy/az

OpenStudy (anonymous):

If you can get your hands on a functions text book the first chapter in those books usually go over all the math rules. WHich they should really make you memorize in grade school but meh

OpenStudy (anonymous):

well basic algebra rules

OpenStudy (anonymous):

learning them will prevent a lot of frustration lol

OpenStudy (anonymous):

:-))) if i told u im way past that part of school u'd laugh :-) i am doing someone a favor and helping HIM with his math ... lol :-) i happen to love math but dont remember it all in short the ans i think should be 1/x^14

OpenStudy (anonymous):

but thank you soooo much!!!!!

OpenStudy (anonymous):

Now I'm doing calc without really studying and still pulling good grades

OpenStudy (anonymous):

:-) no shizz me too! lol im taking this aleks class now for college :-))

OpenStudy (anonymous):

I mean x^(9)/x^(5)

OpenStudy (anonymous):

thats what i got but then i flipped it and multiplied it

OpenStudy (anonymous):

oh you are right lol crud my brian sucks right now been studying really hard

OpenStudy (anonymous):

forgot you can subtract them because they are the same variable

OpenStudy (anonymous):

but yeah good luck

OpenStudy (anonymous):

thanks so much!! peace out man